# Adaptive Requesting in Decentralized Edge Networks via Non-Stationary Bandits

### v2

@@ Abstract
We study a decentralized collaborative requesting problem that aims to optimize the information freshness of time-sensitive clients in edge networks consisting of multiple clients, access nodes (ANs), and servers. Clients request content through ANs acting as gateways, without observing AN states or the actions of other clients. We define the reward as the age of information reduction resulting from a client’s selection of an AN, and formulate the problem as a non-stationary multi-armed bandit. In this decentralized and partially observable setting, the resulting reward process is history-dependent and coupled across clients, and exhibits both abrupt and gradual changes in expected rewards, rendering classical bandit-based approaches ineffective. To address these challenges, we propose the Aging Bandit with Adaptive Reset algorithm, which combines adaptive windowing with periodic monitoring to track evolving reward distributions. We establish theoretical performance guarantees showing that the proposed algorithm achieves near-optimal performance, and we validate the theoretical results through simulations.
@@ I Introduction
To overcome these limitations, modern network designs are increasingly shifting toward decentralized edge networks [29] that distribute computation, storage, and control access the network [9]. Such networks typically consist of end users, access nodes (ANs), and servers, where ANs are deployed closer to end users and serve as intermediate network entities. Within this framework, ANs function as gateways that perform localized caching and forwarding of information, enabling data to be processed and delivered without always traversing the entire network to the cloud. By shortening communication paths and alleviating backhaul congestion, this decentralized network effectively reduces end-to-end latency and allows time-critical updates to be delivered to clients in a more timely manner [9].
@@
Given the challenges above, traditional theoretical analysis becomes infeasible, motivating the use of reinforcement learning techniques, which have demonstrated strong performance in real-time decision-making tasks [41]. In decentralized edge networks, it is essential to develop decentralized learning that allow individual nodes—each with access only to local observations—to collaboratively optimize global objectives without centralized coordination [13]. Among such approaches, Multi-Armed Bandit (MAB) methods [34] are particularly promising due to their simplicity, scalability, and favorable analytical properties, making them well-suited for distributed online decision-making under uncertainty.
@@ I-A Contributions
In this work, we study the problem of optimizing information freshness for time-critical clients in decentralized edge networks. Each client independently selects an AN based solely on its local observations, without access to the states of ANs or the actions of other clients. By defining the reward as the instantaneous reduction in age, we formulate the coordinated request optimization problem as a decentralized Aging Bandit problem with highly non-stationary and correlated rewards, featuring both abrupt and gradual environmental changes.
@@
    Theoretical Guarantees. We develop a theoretical framework for non-stationary multi-armed bandits that accommodates environments in which abrupt and gradual changes coexist. The existing framework (e.g., ADR [23]) typically relies on simplified assumptions that the environment exhibits either purely abrupt or purely gradual changes. In this work, we systematically extend such framework to a more general non-stationary setting with mixed change dynamics. Based on the proposed theoretical framework, we prove that the proposed algorithm is asymptotically optimal, achieving sub-linear regret over time. Extensive simulations further corroborate the theoretical findings.
@@ I-B Notation
The rest of the paper is organized as follows. Section II introduces the system model and problem formulation. Section III defines the age-based reward and reformulates the original problem as a non-stationary multi-armed bandit framework. Section IV presents the proposed ABAR algorithm. Theoretical guarantees are established in Section V and Section VI. Simulation results are reported in Section VII. Finally, Section VIII concludes the paper.
@@ III-B Myopic Reformulation and Reward Design
If client $j$ does not request content from server $p$ via any AN, i.e., $\sum_{k\in[K]}\beta_{jp}^{(k)}(t)=0$, then $x_{jp}(t)=0$.
@@ III-C Non-Stationary MAB and AoI Regrets
The optimization in (III-B) naturally aligns with decentralized online learning and lends itself to a MAB formulation. In particular, for each client-server pair $(j,p)$, the ANs can be viewed as arms, and each client independently interacts with the environment to balance exploration (discovering better ANs) and exploitation (selecting known to yield higher rewards). However, while MAB provides a lightweight framework for decentralized online learning, classical MAB models still remain invalid for our system.
The first reason is the dependence among rewards observed by different clients. Unlike collision-based decentralized MAB models [33, 25], where multiple agents selecting the same arm independently receive zero rewards, our system is fundamentally different. Specifically, when multiple clients simultaneously request content from the same server $p$ via a common AN, the AN makes a single update decision—either fetching a fresh packet from the server or serving cached content. This single update decision affects all requesting clients and further determines their received rewards. As a result, the rewards observed by different clients selecting the same AN are no longer independently. This structural dependence violates the reward independence assumptions commonly adopted in classical multi-armed bandit models.
@@ IV-C Complete Algorithm Description
Compared with the ADR framework [23], ABAR introduces two key extensions. First, the existing ADR framework is built on simplied models that assume changes are either strictly abrupt or strictly gradual. In contrast, in our setting, reward dynamics are history-dependent and evolve through a combination of abrupt and gradual changes. By integrating periodic monitoring with adaptive resets, ABAR does not require prior assumptions about change patterns, enabling reliable adaptation to complex, real-world network dynamics. Second, ABAR operates in a decentralized multi-agent setting, where each client runs its own instance of the algorithm. While clients act independently, shared observations introduce statistical coupling among agents.
@@
5:    $I_{jp}(t)=i_{jp}^{(l-1)}$ (monitor previous block)
@@
7:    $I_{jp}(t)=i_{jp}^{(l)}$ (monitor current block)
@@ V Preliminaries: Notations, Definitions, and Assumptions
We begin by extending the definitions of gradual and abrupt reward changes introduced in [23]. Let $\mu_{i,t}$ denote the expected reward of arm $i$ at time slot $t$.
@@
For any $0\leqslant m\leqslant M$, define
@@
For each $m\in[K]$, assume that
@@ VI Asymptotic Optimality
Combining Theorem 1 with Theorem 2, we obtain that the regret of our algorithm grows sublinearly with the horizon $T$, which implies the algorithm is asymptotically optimal.
@@ VII-A Simulation Setup and Parameter Configuration
We configure the simulation parameters as follows. The time horizon spans $T=6\times 10^{5}$ time slots. The network consists of $J=2$ clients, $K=3$ ANs, and $P=1$ server. Each client sends exactly one request per time slot, i.e., $\sum_{k\in[K]}b_{jp}^{(k)}(t)=1$. To evaluate algorithm robustness under varying network conditions, we consider two different sets of probabilities that the ANs fetch a fresh packet from the server:
@@ VIII Conclusion
To address these challenges, we propose the ABAR algorithm. By combining adaptive windowing with periodic monitoring, ABAR effectively detect changes in reward distributions and promptly discards outdated observations through resent operations. Compared with existing ADR-based framework, ABAR extends the theoretical framework to more general non-stationary setting with mixed change dynamics. We further establish theoretical performance guarantees for ABAR and validate its effectiveness through extensive simulations.

### v3

@@ Abstract
We study a decentralized collaborative requesting problem that aims to optimize the information freshness of time-sensitive clients in edge networks consisting of multiple clients, access nodes (ANs), and servers. Clients request content through ANs acting as gateways, without observing AN states or the actions of other clients. We define the reward as the age of information reduction resulting from a client’s selection of an AN, and formulate the problem as a non-stationary multi-armed bandit. In this decentralized and partially observable setting, the resulting reward process is history-dependent and coupled across clients, and exhibits both abrupt and gradual changes in expected rewards, rendering classical bandits ineffective. To address these challenges, we propose the Aging Bandit with Adaptive Reset algorithm, which combines adaptive windowing with periodic monitoring to track evolving reward distributions. We establish theoretical performance guarantees showing that the proposed algorithm achieves near-optimal performance, and we validate the theoretical results through simulations.
@@ I Introduction
To overcome these limitations, modern network designs are increasingly shifting toward decentralized edge networks [29] that distribute computation, storage, and control across the network [9]. Such networks typically consist of end users, access nodes (ANs), and servers, where ANs are deployed closer to end users and serve as intermediate network entities. Within this framework, ANs function as gateways that perform localized caching and forwarding of information, enabling data to be processed and delivered without always traversing the entire network to the cloud. By shortening communication paths and alleviating backhaul congestion, this decentralized network effectively reduces end-to-end latency and allows time-critical updates to be delivered to clients in a more timely manner [9].
@@
Given the challenges above, traditional theoretical analysis becomes infeasible, motivating the use of reinforcement learning techniques, which have demonstrated strong performance in real-time decision-making tasks [41]. In decentralized edge networks, it is essential to develop decentralized learning that allows individual nodes—each with access only to local observations—to collaboratively optimize global objectives without centralized coordination [13]. Among such approaches, Multi-Armed Bandit (MAB) methods [34] are particularly promising due to their simplicity, scalability, and favorable analytical properties, making them well-suited for distributed online decision-making under uncertainty.
@@ I-A Contributions
In this work, we study the problem of optimizing information freshness for time-critical clients in decentralized edge networks. Each client independently selects an AN based solely on its local observations, without access to the states of ANs or the actions of other clients. By defining the reward as the instantaneous reduction in AoI, we formulate the coordinated request optimization problem as a decentralized Aging Bandit problem with highly non-stationary and correlated rewards, featuring both abrupt and gradual environmental changes.
@@
    Theoretical Guarantees. We develop a theoretical framework for non-stationary multi-armed bandits that accommodates environments in which abrupt and gradual changes coexist. The existing framework (e.g., ADR [23]) typically relies on simplified assumptions that the environment exhibits either purely abrupt or purely gradual changes. In this work, we systematically extend this framework to a more general non-stationary setting with mixed change dynamics. Based on the proposed theoretical framework, we prove that the proposed algorithm is asymptotically optimal, achieving sub-linear regret over time. Extensive simulations further corroborate the theoretical findings.
@@ I-B Notation
The rest of the paper is organized as follows. Section II introduces the system model and problem formulation. Section III defines the AoI-based reward and reformulates the original problem as a non-stationary multi-armed bandit framework. Section IV presents the proposed ABAR algorithm. Theoretical guarantees are established in Section V and Section VI. Simulation results are reported in Section VII. Finally, Section VIII concludes the paper.
@@ III-B Myopic Reformulation and Reward Design
If client $j$ does not request content from server $p$ via any AN, i.e., $\sum_{k\in[K]}\beta_{jp}^{(k)}(t)=0$, then $x_{jp}^{(k)}(t)=0,\,\,\forall\,k\in[K]$.
@@ III-C Non-Stationary MAB and AoI Regrets
The optimization in (III-B) naturally aligns with decentralized online learning and lends itself to a MAB formulation. In particular, for each client-server pair $(j,p)$, the ANs can be viewed as arms, and each client independently interacts with the environment to balance exploration (discovering better ANs) and exploitation (selecting known to yield higher rewards). While MAB provides a lightweight framework for decentralized online learning, classical MAB models still remain invalid for our system.
The first reason is the dependence among rewards observed by different clients. Unlike collision-based decentralized MAB models [33, 25], where multiple agents selecting the same arm independently receive zero rewards, our system is fundamentally different. Specifically, when multiple clients simultaneously request content from the same server $p$ via a common AN, the AN makes a single update decision—either fetching a fresh packet from the server or serving cached content. This single update decision affects all requesting clients and further determines their received rewards. As a result, the rewards observed by different clients selecting the same AN are no longer independent. This structural dependence violates the reward independence assumptions commonly adopted in classical multi-armed bandit models.
@@ IV-C Complete Algorithm Description
Compared with the ADR framework [23], ABAR introduces two key extensions. First, the existing ADR framework is built on simplied models that assume changes are either strictly abrupt or strictly gradual. In contrast, in our setting, reward dynamics are history-dependent and evolve through a combination of abrupt and gradual changes. By integrating periodic monitoring with adaptive resets, ABAR does not require prior assumptions about change patterns, enabling reliable adaptation to complex, real-world network dynamics. Second, ABAR operates in a decentralized multi-agent setting, where each client runs its own instance of the algorithm. While clients act independently, shared observations introduce statistical coupling among agents. By resetting and discarding outdated statistics, ABAR alleviates the impact of such coupling and improves adaptability to non-stationary environments.
@@
5:    $I_{jp}(t)=i_{jp}^{(l-1)}$ (monitoring AN of previous block)
@@
7:    $I_{jp}(t)=i_{jp}^{(l)}$ (monitoring AN of current block)
@@ V Preliminaries: Notations, Definitions, and Assumptions
We begin by extending the definitions of gradual and abrupt reward changes introduced in [23]. Let $\mu_{i,t}$ denote the expected reward of arm $i$ at time slot $t$. Moreover, we denote $i^{(l)}$ as the monitoring arm selected by the algorithm in block $l$.
@@
For any $1\leqslant m\leqslant M$, define
@@
For each $m\in\{1,2,...,M\}$, assume that
@@ VI Asymptotic Optimality
Combining Theorem 1 with Theorem 2, we obtain that the regret of our algorithm grows sublinearly with the time horizon $T$, which implies the algorithm is asymptotically optimal.
@@ VII-A Simulation Setup and Parameter Configuration
We configure the simulation parameters as follows. The time horizon spans $T=6\times 10^{5}$ time slots. The network consists of $J=2$ clients, $K=3$ ANs, and $P=1$ server. Each client sends exactly one request per time slot, i.e., $\sum_{k\in[K]}b_{jp}^{(k)}(t)=1,\,\,\forall\,j\in[J],\,p\in[P],\,t\in[T]$. To evaluate algorithm robustness under varying network conditions, we consider two different sets of probabilities that the ANs fetch a fresh packet from the server:
@@ VIII Conclusion
To address these challenges, we propose the ABAR algorithm. By combining adaptive windowing with periodic monitoring, ABAR effectively detect changes in reward distributions and promptly discards outdated observations through reset operations. Compared with existing ADR-based framework, ABAR extends the theoretical framework to more general non-stationary setting with mixed change dynamics. We further establish theoretical performance guarantees for ABAR and validate its effectiveness through extensive simulations.